MainPage:Nuclear:KaonDetector:PMTCharacteristics:LogBook:pag05
Jump to navigation
Jump to search
⇐ Back to the PMT Characteristics page |
⇒ Go to Summer 2012 work page |
LogBook on the CODA data acquisition, at CUA
Page (last), [6], 5, [4], [3], [2], [1].
Run Number | Date | User | PMT | High Voltage | Experiment | Description and observations | Image |
---|---|---|---|---|---|---|---|
New HV module (VME CAEN V6533N) installed. Scan setup rechecked for bad contacts. Restarting scan project... | |||||||
1224 | 10/01/2013 | Marco | Blast | -1.800 (VME) | Gain | First test to analyze calibration. LED flashing with very low intensity at 1.075kHz. No SEP can be seen. | Error creating thumbnail: Unable to save thumbnail to destination
|
1226 | 10/02/2013 | Marco | Blast | -1.800 (VME) | Gain | Light intensity turned to zero (see only pedestal). | File:1226.png |
1227 | 10/02/2013 | Marco | Blast | -1.800 (VME) | Gain | Low light intensity, but still too much. | Error creating thumbnail: Unable to save thumbnail to destination
|
1228 | 10/02/2013 | Marco | Blast | -1.800 (VME) | Gain | Even lower light intensity. | Error creating thumbnail: Unable to save thumbnail to destination
|
1231 | 10/07/2013 | Marco | Blast | -1.800 (VME) | Scan | Testing LED intensity to start scan run (LED in the center of PMT). | File:1231.png |
1232 | 10/07/2013 | Marco | Blast | -1.800 (VME) | Scan | NO MU-METAL! Scan PMT with LED with pulse width 5 and intensity 7. We can conclude that (a) the PMT is not centralized, (b) the scan device got stuck at some point and (c) I forgot to put Mu-Metal shielding. | Error creating thumbnail: Unable to save thumbnail to destination
|
1233 | 10/08/2013 | Marco | Blast | -1.800 (VME) | Gain | Run to recheck PMT gain after scan. | File:1233.png |
1234 | 10/09/2013 | Marco | Blast | -1.800 (VME) | Scan | Light intensity reajusted. Checking if it was okay, in the center of PMT. Changed a little light intensity after this check. | File:1234.png |
1238 | 10/09/2013 | Marco | Blast | -1.800 (VME) | Scan | NO MU-METAL. Scan PMT with LED with pulse width 6 and intensity 7. No Mu-Metal, just checking if Scan device will get stuck again. Scan device still getting stuck. I changed the final x stop sensor, to see if it solves. Also put more oil on fuse. | Error creating thumbnail: Unable to save thumbnail to destination
|
1240 | 10/10/2013 | Marco | Blast XP4500/B SN09620 | -1.800 (VME) | Gain | MU-METAL INSERTED! Running calib again with LED in the center of PMT. 100ns of trigger window (from NIM quad-four fold logic unit); 6.5 pot pulse width, 6.0 pot pulse intensity (+5.0V). | Error creating thumbnail: Unable to save thumbnail to destination
|
1241 | 10/10/2013 | Marco | Blast XP4500/B SN09620 | -1.000 (VME) | Gain | Calib series, same config as run #1240. | File:1241.png |
1242 | 10/10/2013 | Marco | Blast XP4500/B SN09620 | -1.100 (VME) | Gain | Calib series, same config as run #1240. | File:1242.png |
1243 | 10/10/2013 | Marco | Blast XP4500/B SN09620 | -1.200 (VME) | Gain | Calib series, same config as run #1240. | File:1243.png |
1244 | 10/10/2013 | Marco | Blast XP4500/B SN09620 | -1.300 (VME) | Gain | Calib series, same config as run #1240. | File:1244.png |
1245 | 10/10/2013 | Marco | Blast XP4500/B SN09620 | -1.400 (VME) | Gain | Calib series, same config as run #1240. | File:1245.png |
1246 | 10/10/2013 | Marco | Blast XP4500/B SN09620 | -1.500 (VME) | Gain | Calib series, same config as run #1240. | File:1246.png |
1247 | 10/10/2013 | Marco | Blast XP4500/B SN09620 | -1.600 (VME) | Gain | Calib series, same config as run #1240. | File:1247.png |
1248 | 10/10/2013 | Marco | Blast XP4500/B SN09620 | -1.700 (VME) | Gain | Calib series, same config as run #1240. | File:1248.png |
1249 | 10/10/2013 | Marco | Blast XP4500/B SN09620 | -1.800 (VME) | Gain | Calib series, same config as run #1240. | Error creating thumbnail: Unable to save thumbnail to destination
|
1250 | 10/10/2013 | Marco | Blast XP4500/B SN09620 | -1.900 (VME) | Gain | Calib series, same config as run #1240. | File:1250.png |
1251 | 10/10/2013 | Marco | Blast XP4500/B SN09620 | -2.000 (VME) | Gain | Calib series, same config as run #1240. | File:1251.png |
1252 | 10/10/2013 | Marco | Blast XP4500/B SN09620 | -2.100 (VME) | Gain | Calib series, same config as run #1240. | File:1252.png |
1254 | 10/10/2013 | Marco | Blast XP4500/B SN09620 | -1.800 (VME) | Gain | LED connection changed. Checking calibration. | Error creating thumbnail: Unable to save thumbnail to destination
|
1264 | 10/10/2013 | Marco | Blast XP4500/B SN09620 | -1.800 (VME) | Scan | Checked LED intensity with LED in the center of PMT, to start a scan with high number of PE. LED pot width = 6.0 and intensity = 7.1. | File:1264.png |
1266 | 10/10/2013 | Marco | Blast XP4500/B SN09620 | -1.800 (VME) | Scan | Same conditions as run #1264, starting a scan. | File:1266.png |
1271 | 12/05/2013 | Marco | HMS (SN?) | -1.800 (VME) | Scan | Test scan stability (10x10 scan). Removed T from the cable of LED pulser. | Error creating thumbnail: Unable to save thumbnail to destination
|
1273 | 12/06/2013 | Marco | HMS (SN?) | -1.800 (VME) | Scan | Test scan stability (10x10 scan). Inserted optical fiber at the end of the collimator (end of optical fiber is about 4mm from flat PMT window). | Error creating thumbnail: Unable to save thumbnail to destination
|
1276 | 12/06/2013 | Marco | HMS (SN?) | -1.800 (VME) | Scan | Test scan stability (10x10 scan with 10 points). Removed 2in reference PMT, but black paste around optical fiber and previous hole for reference PMT. It seems like I found the problem: a light leak from the optical fiber (or collimator). All the inserted paste is blocking now the leakage. | Error creating thumbnail: Unable to save thumbnail to destination
|
1277 | 12/06/2013 | Marco | HMS (SN?) | -1.800 (VME) | Scan | Test scan stability (100x100 scan with 30 points in each position). Optical fiber is about 4mm from flat PMT window. | Error creating thumbnail: Unable to save thumbnail to destination
|
1278 | 12/09/2013 | Marco | HMS (SN?) | -1.800 (VME) | Scan | Repeating run 1277, but with a range 2% lower. It is done to test if the end of course is making the horizontal motor to get stucked, the result was a change where it got stucked... | Error creating thumbnail: Unable to save thumbnail to destination
|
1279 | 12/09/2013 | Marco | HMS (SN?) | -1.800 (VME) | Scan | Repeating run 1278, but reducing the motor speed. It was 2600 for both and now it is 1300. Also put some oil on the spindle. This solution didn't fix the problem. | Error creating thumbnail: Unable to save thumbnail to destination
|
1280 | 12/09/2013 | Marco | HMS (SN?) | -1.800 (VME) | Scan | Trying to reduce even more the overall range (4% now). Scan a matrix of 300x300 with 15 points in each position. To avoid overheat, stopping the motors for 1min after each line is scanned. | Error creating thumbnail: Unable to save thumbnail to destination
|
1285 | 12/12/2013 | Marco | HMS (SN?) | -1.800 (VME) | Scan | Came back to 450 step size, 100x100 grid with 30 points in each position. But now moving the vertical motor much more than the horizontal one, so the setup is not getting stuck anymore! | Error creating thumbnail: Unable to save thumbnail to destination
|
1383 | 02/04/2013 | Marco | HMS (SN?) | -1.800 (VME) | Scan | 450 step size, 100x100 grid with 30 points in each position. Some unidentified problem happened. | Error creating thumbnail: Unable to save thumbnail to destination
|
1385 | 02/04/2013 | Marco | HMS (SN?) | -1.600 (VME) | Scan | 450 step size, 100x100 grid with 30 points in each position. Starting the weekly run. | Error creating thumbnail: Unable to save thumbnail to destination
|
1386 | 02/06/2014 | Marco | HMS (SN?) | -1.800 (VME) | Scan | Same as run 1385, just changing HV. | Error creating thumbnail: Unable to save thumbnail to destination
|
1388 | 02/06/2014 | Marco | HMS (SN?) | -2.000 (VME) | Scan | Same as run 1385, just changing HV. | Error creating thumbnail: Unable to save thumbnail to destination
|
1390 | 02/06/2014 | Marco | HMS (SN?) | -1.500 (VME) | Scan | Same as run 1385, just changing HV. | Error creating thumbnail: Unable to save thumbnail to destination
|
1391 | 02/07/2014 | Marco | HMS (SN?) | -1.900 (VME) | Scan | Same as run 1385, just changing HV. | File:1391.png |
1362 | 03/07/2014 | Indra | XP2282B | -2.020 (VME) | Integrating sphere,absorption length of aerogel | Interaction of photon with absorbing material inside the integrating sphere. LED flashing with very low intensity at 1.001kHz. . | File:1462.png |
1369 | 03/07/2014 | Indra | XP2282B | -2.020 (VME) | Integrating sphere,absorption length of aerogel | Interaction of photon with aerogel inside the integrating sphere. LED flashing with very low intensity at 1.001kHz. . | File:1469.png |
1370 | 03/07/2014 | Indra | XP2282B | -2.020 (VME) | Integrating sphere,absorption length of aerogel | Empty integrating sphere. LED flashing with very low intensity at 1.001kHz. . | File:1470.png |
1371 | 03/07/2014 | Indra | XP2282B | -2.020 (VME) | Integrating sphere,absorption length of aerogel | No LED flashing(dark) . | File:1471.png |
1502 | 04/02/2014 | Indra | XP2282B | -1.750 (VME) | Integrating sphere,absorption length of aerogel | LED: on, Aerogel: absent . | File:1502.png |
1504 | 04/02/2014 | Indra | XP2282B | -1.750 (VME) | Integrating sphere,absorption length of aerogel | LED: on, Aerogel: present . | File:1504(1).png |
1505 | 04/02/2014 | Indra | XP2282B | -1.750 (VME) | Integrating sphere,absorption length of aerogel | LED : off, Aerogel present . | File:1505.png |